I need help solving these problems with complex numbers
write the sum as a complex number in standard form
(-5+5i)+(-2-3i)
and
(-3+5i)(-4+3i)

Answer by Prithwis(166)

(-5+5i)+(-2-3i) = (5i-3i)-(5+2) = 2i-7
------------------------------------------------------
(-3+5i)(-4+3i) = 15i^2-20i-9i+12 = -15-29i+12 = -29i-3